OpenStudy (anonymous):

Let u = <-4, 1>, v = <-1, 6>. Find -2u + 4v.

OpenStudy (anonymous):

@agent0smith

jimthompson5910 (jim_thompson5910):

u = <-4, 1> is given to find -2u, you multiply everything in vector u by -2 so, u = <-4, 1> -2u = -2<-4, 1> -2u = <-2(-4), -2(1)> -2u = <8, -2>

jimthompson5910 (jim_thompson5910):

If v = <-1, 6>, then 4v = ???

OpenStudy (anonymous):

<4, 22> <4, 7> <12, -26> <10, -14>

jimthompson5910 (jim_thompson5910):

Do you see how I got -2u?

OpenStudy (anonymous):

so it would be <4, 22> ? @jim_thompson5910

OpenStudy (anonymous):

wait no

jimthompson5910 (jim_thompson5910):

If v = <-1, 6>, then 4v = ???

OpenStudy (anonymous):

<12, -26>

jimthompson5910 (jim_thompson5910):

to calculate 4v, multiply everything in v by 4

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

@jim_thompson5910 then i added them and got <12, -26>

jimthompson5910 (jim_thompson5910):

tell me what you got for 4v

OpenStudy (anonymous):

-4,24

OpenStudy (anonymous):

4,22

jimthompson5910 (jim_thompson5910):

so -2u + 4v turns into <8, -2> + <-4,24>

jimthompson5910 (jim_thompson5910):

Then you add up the corresponding components 8 + (-4) = 4 -2 + 24 = 22 So yes, it's <4, 22>

OpenStudy (anonymous):

thank you for explaining c:

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

@jim_thompson5910 do you know how to explain this by any chance? Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 4>, v = <9, -9>

jimthompson5910 (jim_thompson5910):

are you familiar with dot products?

OpenStudy (anonymous):

no :(

jimthompson5910 (jim_thompson5910):

it turns out that if u = <a,b> v = <c,d> then u dot v = a*c + b*d

jimthompson5910 (jim_thompson5910):

so you multiply the corresponding entries, then add up those products

jimthompson5910 (jim_thompson5910):

if u = <8, 4>, v = <9, -9>, then u dot v = ??

OpenStudy (anonymous):

8*9 + 4*-9 72+ -36

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

36

OpenStudy (anonymous):

but my answer choices are : 81.6° 25.8° 35.8° 71.6° so would it be 35.8? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok so that's part of finding the angle

jimthompson5910 (jim_thompson5910):

idk, we haven't finished yet

jimthompson5910 (jim_thompson5910):

now we must find the lengths of vector u and vector v

OpenStudy (anonymous):

oh okay sorry c:

jimthompson5910 (jim_thompson5910):

if u = <8, 4>, then |u| = ??

jimthompson5910 (jim_thompson5910):

|u| is the length of vector u

OpenStudy (anonymous):

-8 -4?

jimthompson5910 (jim_thompson5910):

|u| = sqrt( u dot u )

jimthompson5910 (jim_thompson5910):

so you first compute the dot product u dot u then take the square root

OpenStudy (anonymous):

wow you lost me.. im so sorry i really suck at this

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

u is a vector, so it has a length and direction

jimthompson5910 (jim_thompson5910):

the length of vector u is found by first computing the dot product of u dot u what is that dot product?

jimthompson5910 (jim_thompson5910):

I'm guessing you're stuck on how to find u dot u?

OpenStudy (anonymous):

yes,,

jimthompson5910 (jim_thompson5910):

did you see how to find u dot v?

OpenStudy (anonymous):

thats the one that i got 36 for right?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you would do the same, but instead of use v, you use vector u

jimthompson5910 (jim_thompson5910):

so you just use the same vector twice

OpenStudy (anonymous):

162?

jimthompson5910 (jim_thompson5910):

I'll show you what I mean u = <8, 4> u dot u = 8*8 + 4*4 u dot u = 64 + 16 u dot u = 80

jimthompson5910 (jim_thompson5910):

or another way to put it <8, 4> dot <8,4> = 8*8 + 4*4 = 80

jimthompson5910 (jim_thompson5910):

So the length of vector u is sqrt(80) units long

jimthompson5910 (jim_thompson5910):

Using this idea, how long is vector v?

OpenStudy (anonymous):

81.6°

jimthompson5910 (jim_thompson5910):

how long is vector v?

jimthompson5910 (jim_thompson5910):

or, tell me what <9, -9> dot <9, -9> is equal to

OpenStudy (anonymous):

162

jimthompson5910 (jim_thompson5910):

good, so the length of vector v is sqrt(162)

jimthompson5910 (jim_thompson5910):

we now multiply the two lengths to get sqrt(80)*sqrt(162)

jimthompson5910 (jim_thompson5910):

We'll then divide the dot product u dot v by sqrt(80)*sqrt(162) and then finally, take the arccosine of that result

jimthompson5910 (jim_thompson5910):

so 36/( sqrt(80)*sqrt(162) ) then the arccosine of that

OpenStudy (anonymous):

0.31622776601

jimthompson5910 (jim_thompson5910):

now take the arccosine of that

jimthompson5910 (jim_thompson5910):

make sure you are in degree mode

OpenStudy (anonymous):

1.24904577

jimthompson5910 (jim_thompson5910):

seems like you're not in degree mode

OpenStudy (anonymous):

i dont really have a calculator so im using googles

jimthompson5910 (jim_thompson5910):

ok, then you would type in "arccos(0.31622776601) in degrees" type that without quotes

jimthompson5910 (jim_thompson5910):

you have to stick on "in degrees" because google uses radians by default

OpenStudy (anonymous):

71.6

jimthompson5910 (jim_thompson5910):

what decimal number did you get (before you rounded)

OpenStudy (anonymous):

71.5650512

jimthompson5910 (jim_thompson5910):

very good, so 71.6 is correct

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

to sum things up, the angle theta is found by using the formula \[\Large \theta = \arccos\left(\frac{u \cdot v}{|u||v|}\right)\]